OpenStudy (anonymous):

Anyone wanna help? A point T on a segment with endpoints D (1, 4) and F (7, 1) partitions the segment in a 2:1 ratio. Find T.

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

find distance by x: 7-1=6. now multiply 6*2/3=4. Now find y diastnace: 3. 3*2/3=2. 4, 2 now add (1+4) and subtarct (4-2). then T=(5, 2)

OpenStudy (anonymous):

Thankss!! @AlexandervonHumboldt2

OpenStudy (jdoe0001):

hmmm if the point T splits the line DF in a 2:1 ratio how many segments is it splitting it in?

OpenStudy (jdoe0001):

hint: 2 to 1 ratio